OpenStudy (anonymous):

Tania bought 4 more pounds of pears then Wilma. Together, Tania and Wilma bought 18 pounds of pears. How many pounds of pears did Wilma buy? Need help now!!!

OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well let Wilma buy x pounds.... then Tania bought x + 4 pounds... so together you get the equation x + x+ 4 = 18 or 2x + 4 = 18 if you solve for x you can find what each person purchased.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

kill yourself

OpenStudy (anonymous):

24? @campbell_st

OpenStudy (campbell_st):

nope... how can the answer be 24 when the combined weight purchased was only 18 pounds.

OpenStudy (anonymous):

wait no! that was dumb!

OpenStudy (anonymous):

5? @campbell_st

OpenStudy (campbell_st):

nope... here is what you have 2x + 4 = 18 subtract 4 from both sides - 4 -4 ----------- 2x = 14 so what how many pounds did Wilma buy...?